OpenStudy (anonymous):

The ordered pairs (1,4) (2,16) (3,64) and (4, 256) represent a function. What is a rule that represents this function? a) y = x^4 b) y = 4^x c) y = 4x d) y = x + 4

OpenStudy (usarmy3947):

let me see how i can help u : ) @abb0t @dan815

OpenStudy (anonymous):

the answer is b)y = 4^x

OpenStudy (anonymous):

let's see what we have here\[4=4^1\]\[16=4^2\]\[64=4^3\]...